OpenStudy (anonymous):

Hey guys if you can help me i would really appreciate it! F(x) = __3x+4___ 5x+9 The domain of (f) contains all real numbers x , except x= ???????????

OpenStudy (amistre64):

is this a division problem?

OpenStudy (amistre64):

if it is, then the law states that we cannot under any circumstances divide by zero. It will rip a hole the size of nantucket into the very fabric of reality into which nothing can escape.... what would make the denominator, the bottom, equal zero?

OpenStudy (anonymous):

haha ok, and yes it is!

OpenStudy (anonymous):

would we multiply by zero

OpenStudy (amistre64):

nope.... we would do this: 5x+9=0 5x +9-9 = 0-9 5x +0 = -9 5x = -9 5x/5 = -9/5 1x = -9/5 x cannot under any circumstances be allowed be equal to -9/5. So keep it out of the cupboards and make sure you dont feed it after midnight.....